OpenStudy (anonymous):

log(lower 3)((x^2)+8)=2 need the solution set PLEASE help im doing a timed homework assignment

OpenStudy (anonymous):

Do you remember what a log base anything is?

OpenStudy (anonymous):

noooo :((

OpenStudy (anonymous):

hmm Okay... Do you remember how to turn a log into exponential form?

OpenStudy (anonymous):

its 1 and -1

OpenStudy (anonymous):

i answered this somewhere else?

OpenStudy (anonymous):

beacuse its 3 ^2 = x^2 +8...which means x^2 = 9-8, which equals 1., so x can be either 1 or -1

OpenStudy (anonymous):

ohhhh ok! thank you!!!